OpenStudy (anonymous):

-3/2=4/5w+1/4 what is w? please help me :) i award

OpenStudy (anonymous):

w=-35/16

OpenStudy (anonymous):

=-2.1875

OpenStudy (anonymous):

6/5u-4/3=-5/6?

OpenStudy (anonymous):

u=5/12

OpenStudy (anonymous):

≈0.4166666666667

OpenStudy (anonymous):

1 more!! -5/6+7/8w=-3/2

OpenStudy (anonymous):

w=-16/21 ≈-0.7619047619048}

OpenStudy (anonymous):

Thank you so so so much!!

OpenStudy (anonymous):

lol oks!!